### v1

# Efficient State Estimation of a Networked FlipIt Model

Abstract
The Boolean Kalman Filter and associated Boolean Dynamical System Theory have been proposed to study the spread of infection on computer networks. Such models feature a network where attacks propagate through, an intrusion detection system that provides noisy signals of the true state of the network, and the capability of the defender to clean a subset of computers at any time. The Boolean Kalman Filter has been used to solve the optimal estimation problem, by estimating the hidden true state given the attack-defense dynamics and noisy observations. However, this algorithm is infeasible because it runs in exponential time and space with respect to the network size. We address this feasibility problem by proposing a mean-field estimation approach, which is inspired by the epidemic modeling literature. Although our approach is heuristic, we prove that our estimator exactly matches the optimal estimator in certain non-trivial cases. We conclude by using simulations to show both the run-time improvement and estimation accuracy of our approach.
I Introduction
Cybersecurity has been extensively studied with mathematical treatments while assuming full and accurate information (see, e.g. [1, 2]). However, the StuxNet incident [3] served as a “wake-up call” on studying partial and noisy information, dubbed low-information, where partial means defenders’ sensors are only employed at limited places and noisy means the sensors provide inaccurate information. This can be attributed to both the inadequacy of sensor employments (mandated by cost constraints) and the sophistication of attacks, especially the new or zero-da attacks waged by nation-state actors, also known as Advanced Persistent Threats (APTs).
For modeling APTs, one family of studies are centered around the FlipIt model [4]. The original FlipIt model is built around three core assumptions. (i) There is a computer that an attacker and defender can seize control over it at any time. This is often interpreted as the attacker using an APT to gain control and the defender reinstalling the OS or reimaging a virtual machine. (ii) Neither the attacker nor the defender has any information about who currently controls the computer. (iii) Seizing control of the computer comes at a cost, in terms of effort by the attacker and the lost uptime to the defender. Thus, FlipIt can be seen as a low-information model to study how often to deploy security measures against a capable attacker.
Although many extensions to FlipIt have been proposed, including multiple computers [5] and insider threats [6]. This family of models has two weaknesses. (i) They consider an individual computer, rather than networks. In the real world, computers are networked, resulting in much more complicated scenarios as evidenced by cascading effects and lateral movements (after penetrating into a network). (ii) The low-information assumption does not reflect reality in many settings because Intrusion Detection Systems (IDSes) are widely deployed both on computers (or hosts) and in networks to detect cyber attacks. Although far from perfect, IDSes provide a noisy reading of the true states of computers, which is not modeled by FlipIt-style models.
Recently, the theory of Boolean Dynamical Systems [7] has been applied to study the APT problem with both noisy IDSes and network structures. Originally proposed as a discrete-time dynamical system where each node has a state 00 or $1$ and the dynamics are described by arbitrary modulo 2 operations between connected nodes. The network structures and noisy observations introduce the following non-trivial state estimation problem (in addition to the FlipIt-style decision problem): What is the probability that a given computer is compromised (or infected) at a specific point in time given the network structure and the history of IDS signals (i.e., alerts)? Although this estimation problem has been optimally solved by the Boolean Kalman Filter (BKF) [8], it incurs an exponential time and space in terms of the size of the network because it requires computation over every possible combination of network-wide computer states.
In this work, we propose a novel estimator, dubbed the Mean Field Analysis (MFA) estimator, because its equations resemble the MFA dynamics from networked epidemic spreading models [9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. In the proposed MFA method, we keep track of the probability of infection for each node rather than a distribution over all possible states, leading to a polynomial-time estimation algorithm. Our results are summarized as follows:
  1. 1.
    We propose a novel estimator (the MFA algorithm) based on an approximation of Bayes’ Law using the Maximum Entropy Principle.
  2. 2.
    We prove that all steps of the MFA algorithm are exact and efficient computations of the corresponding step of the BKF algorithm, under the condition that the prior belief is an appropriate maximum entropy distribution (Theorems 2,3).
  3. 3.
    We run simulations to compare the performance of the MFA and BKF estimators.
Thus, we provide a novel estimator that is both theoretically justified and practical for running on large real-world networks. Notably, our proposed method is suitable for real-time computation on large networks with hundreds of nodes, where the state-of-the-art approach BKF takes hours on graphs with just 14 nodes.
II Problem Formulation
II-A Model
We model the spread of cyber attacks in a computer network where the true state of each node is unknown. This can be considered a hidden Markov model with boolean states and given noisy state observations at each discrete time step. We consider networks with $n$ nodes and let $\{x_{t};t=0,1,2,\dots\}$ be a state process where $x_{t}\in\{0,1\}^{n}$ is the state of the process at time $t$. We use $x_{t,l}$ to denote the state of computer $l$ at time $t$ where $x_{t,l}=1$ means the computer is compromised and $x_{t,l}=0$ means the computer is secure. Let $X=[x^{1},x^{2},\ldots,x^{2^{n}}]$ be an $n\times 2^{n}$ matrix where each $x^{i}\in\{0,1\}^{n}$ represents a unique state. Together, all of the columns $x^{i}$ of $X$ include all possible states in $\{0,1\}^{n}$. Occasionally, it is convenient to abuse notations somewhat and treat $X$ as a set, namely $X=\{x^{1},x^{2},\dots,x^{2^{n}}\}$. We denote a distribution over all $2^{n}$ states as $\Pi\in\Delta(X)$, and $P_{t}\in[0,1]^{n}$ a vector of distributions of each node. To avoid confusion, we use consistent symbols for indexing, $t\in\{0,1,2,\ldots\}$ is a time index, $l,k,r\in V$ are node index, and $i,j\in\{1,2,\ldots,2^{n}\}$ are index over $X$ (e.g., $x^{i},x^{j}\in X$). Due to the amount of indexing required, we only use time indexes, when necessary, to describe the evolution of $x_{t},\Pi_{t}$ and $P_{t}$ over time.
Underlying the computers is a directed graph or network structure $G=(V,E)$ where the set of computers (i.e., nodes) is denoted by $V=\{1,2,\ldots,n\}$, and the links (or connections) between them are given by directed edge set $E$. A computer $l$ can propagate an attack to computer $k$ if $(l,k)\in E$. The set of connections models the communications permitted by the cybersecurity policy (known as attack-defense structure [12]), rather than representing the physical network structure or topology. Let $D_{l}=\{k\in V\mid(k,l)\in E\}$ be the in-neighbor set of a node $l$.
At each time step, if an in-neighbor $k$ of node $l$ is compromised, then the attack will successfully compromise node $l$ with probability $\rho_{kl}\in[0,1]$. At each time step the defender may opt to clean each node individually, indicated by $a_{t}\in\{0,1\}^{n}$, where $a_{t,l}=1$ means the defender cleans node $l$ at time $t$ even if the defender is not certain whether node $l$ is compromised or not, and conversely for $a_{t,l}=0$. For model generality, we assume the cleaning will have a failure probability $\alpha\in[0,1]$, meaning a cleaning will succeed with probability $(1-\alpha)$. Finally, we assume that node $1$ models an external threat and is always compromised, namely $x_{t,1}=1$ for all $t$, and cannot be cleaned, namely $a_{t,1}=0$ for all $t$. Particularly, any node $l$ that has an edge from node 1, namely $(1,l)\in E$, is thought to be vulnerable to external threats and any subset of nodes may have connections from node 1. This assumption is made out of mathematical convenience and can be seen as equivalent to the environment (in which a network resides) such that the nodes in the network may be susceptible to pull-based (e.g., drive-by download) attacks that were first mathematically modeled in [12]); moreover, [19] considers an equivalent version where the attack is not explicitly modeled in the network.
The dynamics of the true network state is given by
$\Pr(x_{t+1,l}=1\mid x_{t},a_{t})$ $=(1+(\alpha-1)a_{t,l})\bigg{(}x_{t,l}+$ (1)
$(1-x_{t,l})\bigg{[}1-\prod_{k\in D_{l}}(1-\rho_{kl}x_{t,k})\bigg{]}\bigg{)},$
which gives the probability that node $l$ is compromised given the previous state and cleaning action $(x_{t},a_{t})$, the graph structure, and the attack success parameters $\rho_{kl}$. The cleaning success term $(1+(\alpha-1)a_{t,l})$ denotes the probability the node is not cleaned successfully, and the subsequent term denotes the probability that $l$ is compromised given the previous state of the network $x_{t}$. Particularly, if $l$ was not previously compromised, namely $x_{t,l}=0$, then the compromise probability is determined to the term with the product, which is the probability at least one successful attack comes from any in-neighbor.
After the attack and cleaning process has concluded, we assume that we receive a noisy measurement of the true state of the node, denoted by $y_{t}\in\{0,1\}^{n}$, where $y_{t,l}=1$ indicates that IDS (potentially incorrectly) detected $x_{t,l}=1$ and likewise for $y_{t,l}=0$, namely that the noises are incurred by the False Positives and False Negatives of IDS. The observation $y_{t}$ is noisy as follows
$y_{t,l}=\begin{cases}1\mbox{ if }x_{t,l}=1&\mbox{with probability }p\\ 0\mbox{ if }x_{t,l}=1&\mbox{with probability }1-p\\ 0\mbox{ if }x_{t,l}=0&\mbox{with probability }q\\ 1\mbox{ if }x_{t,l}=0&\mbox{with probability }1-q\\ \end{cases}$ (2)
where $p,q\in[0,1]$ are the True Positive and True Negative probabilities respectively. Occasionally, we use $\tilde{p}=1-p$ and $\tilde{q}=1-q$ for compactness.
Given that the true state $x_{t}$ is unknown but dynamics (1), cleanings $a_{t}$, and observations $y_{t}$ are known, we seek an estimator $\hat{x}_{t}\in\{0,1\}^{n}$ of the true state $x_{t}$. To measure the quality of an estimator, we use the mean-squared error (MSE) criterion
$C(x_{t},\hat{x}_{t})=\mathbb{E}(||x_{t}-\hat{x}_{t}(a_{0:t-1},y_{1:t})||^{2}_{2}\mid a_{0:t-1},y_{1:t})$ (3)
where $||\cdot||_{2}$ is the $\ell_{2}$ norm. This objective function counts the expected number of nodes incorrectly estimated given the history of observations and cleanings. Thus, we seek estimators that can (approximately) achieve optimization with respect to MSE, namely
$\hat{x}^{*}_{t}\in\operatorname*{arg\,min}_{\hat{x}_{t}\in\Psi}C(x_{t},\hat{x}_{t})$ (4)
where $\Psi$ is the set of all functions that map $a_{0:t-1},y_{1:t}$ onto $\{0,1\}^{n}$.
II-B The Boolean Kalman Filter
Data: $\Pi_{0\mid 0}$
Result: $\hat{x}^{*}_{1},\hat{x}^{*}_{2},\hat{x}^{*}_{3},\cdots$
for t=1,2,… do
      1 $\Pi_{t|t-1}\leftarrow M_{t}(a_{t-1})\Pi_{t-1|t-1}$;
      2 $\Pi_{t|t}\leftarrow\frac{T_{t}(y_{t})\circ\Pi_{t\mid t-1}}{(T_{t}(y_{t}))^{\top}\Pi_{t\mid t-1}}$;
      3 $\hat{x}^{*}_{t}=\overline{X\Pi_{t|t}}$;
end for
Algorithm 1 The Boolean Kalman Filter Algorithm
The optimal estimator on the preceding problem, known as the Boolean Kalman Filter (BKF), has been given in [19] (with minor technical differences). Before giving the full details of the algorithm, we discuss the high-level ideas behind it. The algorithm has three steps at each iteration, where the probability distribution $\Pi_{t-1|t-1}\in\Delta(X)$ is updated to $\Pi_{t|t}\in\Delta(X)$, which is leveraged to produce the optimal MSE estimator. The distribution $\Pi_{t-1|t-1}$ is known as the prior belief and is the existing belief across all states $X$. In the first step, the matrix-vector product is taken
$\Pi_{t|t-1}=M_{t}(a_{t-1})\Pi_{t-1|t-1},$ (5)
where $M_{t}$ is a $2^{n}\times 2^{n}$ column stochastic matrix that represents the expected dynamics of (1) with the cleaning vector $a_{t-1}$. This produces an intermediary belief $\Pi_{t|t-1}$, which is then operated on by vector $T_{t}(y_{t})$ to obtain the updated belief $\Pi_{t\mid t}$, as follows:
$\Pi_{t\mid t}=\frac{T_{t}\circ\Pi_{t\mid t-1}}{(T_{t})^{\top}\Pi_{t\mid t-1}},$ (6)
where $\circ$ denotes the element-wise multiplication operator and $\top$ denotes the transpose operator. The vector $T_{t}(y_{t})$ leverages the actual observation $y_{t}$ with the True Positive and True Negative probabilities $p,q$ to adjust the values of belief $\Pi_{t\mid t}$ accordingly. Finally, the optimal MSE estimate is provided by
$\hat{x}_{t}=\overline{X\Pi_{t\mid t}},$ (7)
where the overline operator $\overline{\cdot}$ is the element-wise rounding operator. The drawback of the BKF is that all three components, $\Pi_{t\mid t}$,$M_{t}$,$T_{t}$, are all exponentially sized, namely at the magnitude of $2^{n}$. This automatically implies the BKF algorithm is not feasible for sufficiently large networks.
We now provide the full details of $M_{t}(a_{t-1})$ and $T_{t}(y_{t})$, beginning with $M_{t}(a_{t-1})$. Given state $x^{j}\in X$, the probability that node $l$ is compromised in the next time step is given by
$\eta^{j}_{l}=$ $\bigg{(}1+a_{t-1,l}(\alpha-1)\bigg{)}$ (8)
$\quad\bigg{(}x^{j}_{l}+(1-x^{j}_{l})\bigg{[}1-\prod_{r\in D_{l}}1-\rho_{rl}x^{j}_{r}\bigg{]}\bigg{)}$
which directly follows the dynamics (1), while taking into account cleaning $a_{t-1}$ and the previous state of node $x^{j}_{l}$. We leverage $\eta^{j}_{l}$ to define the column stochastic matrix $M_{t}$, which is the transition matrix that defines the hidden Markov chain over the $2^{n}$ states. Specifically, entry $(M_{t})_{ij}$ is the probability that state $x^{j}$ will transition to $x^{i}$, namely:
$(M_{t})_{ij}$ $=\Pr(x_{t}=x^{i}\mid x_{t-1}=x^{j},a_{t-1})$ (9)
$=\prod^{n}_{l=1}\bigg{(}\eta^{j}_{l}x^{i}_{t,l}+(1-\eta^{j}_{l})(1-x^{i}_{t,l})\bigg{)}.$
The matrix $M_{t}$ fully describes the expected underlying dynamics of the hidden Markov model, but observation $y_{t}$ is available to improve the estimate. To leverage $y_{t}$, we construct vector $T_{t}\in[0,1]^{2^{n}}$ such that
$(T_{t}(y_{t}))_{i}$ $=\Pr(y_{t}\mid x_{t}=x^{i})$ (10)
$=\prod^{n}_{l=1}\Pr(y_{t,l}\mid x_{t}=x^{i})$
$=\prod^{n}_{l=1}\bigg{[}y_{t,l}\bigg{(}px^{i}_{l}+(1-q)(1-x^{i}_{l})\bigg{)}$
$\qquad+(1-y_{t,l})\bigg{(}q(1-x^{i}_{l})+(1-p)x^{i}_{l}\bigg{)}\bigg{]}$
which is the probability that observation $y_{t}$ occurred given that the true state was $x^{i}$.
III Approach
III-A An Approximate Approach
To remedy the computational complexity of the BKF algorithm, we present a polynomial-time approximation based on Bayes’ Law conditioned on the belief that each node is compromised. Particularly, we propose tracking only the belief $P_{t}\in[0,1]^{n}$ that corresponds to the probability that each node is compromised. This corresponds to the product $X\Pi_{t\mid t}$, which has the same interpretation. In particular, we use Bayes’ Law to condition on the fact that we only have belief $P_{t}\in[0,1]^{n}$ as the prior instead of $\Pi_{t\mid t}\in\Delta(X)$. The dynamics can then be given by
$P_{t+1,l}=\Pr(x_{t+1,l}=1\mid P_{t},a_{t},y_{t+1})=$ (11)
$\begin{cases}\frac{\Pr(y_{t+1,l}=1\mid P_{t},a_{t},x_{t+1}=1)\Pr(x_{t+1,l}=1\mid P_{t},a_{t})}{Pr(y_{t+1,l}=1\mid P_{t},a_{t})}&y_{t+1,l}=1\\ \frac{(1-\Pr(y_{t+1}=1\mid P_{t},a_{t},x_{t+1}=1))\Pr(x_{t+1,l}=1\mid P_{t},a_{t})}{(1-Pr(y_{t+1,l}=1\mid P_{t},a_{t}))}&y_{t+1,l}=0\end{cases}$
via Bayes’ Law. It then suffices to derive each of the three terms as a function of $P_{t},u_{t},y_{t+1}$ to complete our derivation of the MFA estimator. The first term of the numerator is easy to show directly, $\Pr(y_{t+1,l}=1\mid P_{t},a_{t},x_{t+1}=1)=p$; then, by using this result and the law of total probability, the denominator can be solved as
$\Pr(y_{t+1,l}$ $=1\mid P_{t},a_{t})=p\Pr(x_{t+1,l}=1\mid P_{t},a_{t})$
$\quad+\tilde{q}(1-\Pr(x_{t+1,l}=1\mid P_{t},a_{t}))$
which is described using the remaining term $\Pr(x_{t+1,l}=1\mid P_{t},a_{t})$. This term is potentially difficult to compute because many distributions $\Pi\in\Delta(X)$ can be consistent with $P_{t}\in[0,1]^{n}$, which we denote by set $S(P_{t})=\{\Pi\in\Delta(X)\mid X\Pi=P_{t}\}$. This raises two questions. First, evaluating $\Pr(x_{t+1,l}=1\mid P_{t},a_{t})$ likely requires finding a distribution over $S(P_{t})$, rather than a single distribution $\Pi$. It is conceivable that novel dynamics would need to be derived because such a distribution would evolve given $y_{t+1},a_{t}$ and dynamics (1). Second, to numerically evaluate that term would require sampling $S(P_{t})$ with the desired distribution, which itself is non-trivial owing to the fact that $X\Pi_{t|t}=P_{t}$ defines a system of $n$ equations that have $2^{n}$ variables.
Regardless, it is likely any computation over the whole set of $S(P_{t})$ would take place in exponential time, which is also true for computing $X\Pi_{t|t}$, leading to the presumption that directly evaluating $\Pr(x_{t+1,l}=1\mid P_{t},a_{t})$ is not tractable.
III-B The Mean Field Analysis Estimator
Data: $P_{0}\in[0,1]^{n}$
Result: $\hat{x}_{1},\hat{x}_{2},\hat{x}_{3},\cdots$
for t=1,2,… do
       for $l\in V$ do
            1 $P^{\prime}_{t,l}\leftarrow(1+(\alpha-1)a_{t,l})$
             $\bigg{(}P_{t-1,l}+(1-P_{t-1,l})\bigg{[}1-\prod_{k\in D_{l}}(1-\rho_{kl}P_{t-1,k})\bigg{]}\bigg{)}$;
            2 $P_{t,l}\leftarrow\frac{py_{t,l}P^{\prime}_{t,l}+(1-y_{t,l})\tilde{p}P^{\prime}_{t,l}}{y_{t,l}[pP^{\prime}_{t,l}+\tilde{q}(1-P^{\prime}_{t,l})]+(1-y_{t,l})[\tilde{p}P^{\prime}_{t,l}+q(1-P^{\prime}_{t,l})]}$;
       end for
[…]
end for
Algorithm 2 The Mean Field Analysis Algorithm
Given that we cannot directly evaluate Bayes’ Law with belief $P_{t}$, the main premise of our estimator is to leverage the Mean Field Analysis (MFA) equation to provide a tractable equation in place of $\Pr(x_{t+1,l}=1\mid P_{t},a_{t})$. The MFA equations have long been proposed for a similar purpose in epidemic modeling (see, e,g,m [9, 15]), where each node is modeled individually. The dynamics are given to (1), with the main difference that in epidemic modeling, there is no notion of cleaning $a_{t}$, but rather some natural recovery process is assumed. However, both models lead to dynamics similar to (5) where there exists an underlying $2^{n}$ state Markov chain that is difficult to analyze. Thus, the so-called MFA equation has been proposed, given in Line 2 of Algorithm 2. Particularly, it can be seen that the MFA equation is structurally similar to dynamics (1) with the exception that states $x_{t}$ have been replaced with belief $P_{t}$. With respect to networked models, there have been considerable numerical efforts to evaluate the performance of Mean Field estimators [16, 18], and it has been shown via two prominent epidemic models [10, 11] that in a special case, the variations of the MFA dynamics do converge to the true dynamics.
We now provide a high-level description of the proposed Algorithm 2that aims to approximate the results of the BKF Algorithm 1. Both algorithms function in 3 sub-steps (Line 1-3) at each iteration; each step of the MFA algorithm functions analogously to the BKF. Note that the MFA algorithm requires an extra loop over $V$ as the matrix products can no longer be used to do that implicitly. Further, it can be seen that Line 2 utilizes the MFA equation to estimate the expected dynamics (1), given prior belief $P_{t-1}$. This produces an intermediate distribution $P^{\prime}_{t}$, which is then updated in Line 2 with respect to the received observation $y_{t}$ utilizing Bayes’ law as given in (11). This update takes into consideration the proposed approximation $\Pr(x_{t,l}=1\mid P_{t-1},a_{t})\approx P^{\prime}_{t}$. Finally, in Line 2, we round the posterior belief $P_{t}$ in an element-wise fashion to obtain the MFA estimator $\hat{x}_{t}$. It should be noted that, at each time step, the algorithm requires an operation over each node $l\in V$, and within that loop, a loop over all in-neighbors of $l$. Thus, the MFA algorithm can be pessimistically regarded as running in $\mathcal{O}(nD^{*})$ time, where $D^{*}=\max_{l\in V}|D_{l}|$ is the maximum number of in-neighbors among all nodes.
IV Results
Although justified in the context of epidemic models, using the MFA equation to evaluate an approximate Bayes’ law is not well justified yet in the context of state estimation. To address this gap, we now prove that all three steps of the MFA algorithm can be derived from the BKF via a special case of $\Pi\in S(P)$. To see this, we revisit the attempt to compute (rather estimate) $\Pr(x_{t+1,l}=1\mid P_{t},a_{t})$. As discussed previously, computing over a distribution over $S(P)$ is difficult and is likely to take exponential time. Thus, one approach to estimating this probability is to select a single $\Pi\in S(P)$, but how to pick? A disciplined answer is the distribution $\Pi\in S(P)$ that maximizes entropy. Formally, given $P\in[0,1]^{n}$, we define
$\Pi^{*}(P)\in\operatorname*{arg\,max}_{\Pi\in S(P)}\sum_{i}-\Pi_{i}\ln\Pi_{i},$ (12)
where $\Pi^{*}(P)$ is the distribution that maximizes the entropy, constrained to be consistent with belief $P$. To justify why the maximum entropy distribution is a disciplined choice, we appeal to the Maximum Entropy Principle. This principle addresses a common problem in statistical estimation, where many possible underlying distributions are consistent with available information. The authors [20] describe the appeal of selecting the maximum entropy distribution as “what we need is a probability distribution which ignores no possibility subject to the relevant constraints.” The intuition is as follows: although it would be more desirable to have a lower entropy distribution, there is no way to select the correct lower entropy distribution consistent with the observations.
Before we present the results on deriving each step of the MFA algorithm, we first give a theorem that derives $\Pi^{*}(P)$ in closed form.
Theorem 1
Given any belief $P\in[0,1]^{n}$, the distribution
$\Pi^{*}(P)\in\Delta(X)$ defined by
$\Pi^{*}(P)_{i}=\prod_{l\in V}(P_{l}x^{i}_{l}+(1-x^{i}_{l})(1-P_{l}))$ (13)
for all $x^{i}\in X$ is the unique distribution that solves (12).
Theorem 1 states that the distribution $\Pi^{*}(P)$ that maximizes entropy while being consistent with the belief $P$ is the product distribution over $X$ induced by $P$. The proof of this result is deferred to the Appendix (Section VI). Using this form, we now evaluate $\Pr(x_{t+1,l}=1\mid\Pi^{*}(P_{t}),a_{t})=XM_{t}\Pi^{*}(P_{t})$, which is Line 1 of the BKF Algorithm 1.
Theorem 2
Let $M_{t}$ be constructed appropriately via (9) for any $a_{t-1}$. Then, $\Pr(x_{t+1,l}=1\mid\Pi^{*}(P_{t}),a_{t})$ may be computed explicitly in $\mathcal{O}(n)$ via
$\Pr(x_{t+1,l}=1\mid\Pi^{*}(P_{t}),a_{t})=(XM_{t}\Pi^{*}(P_{t}))_{l}$ (14)
$=(1+(\alpha-1)a_{t,l})$
$\qquad\left(P_{t,l}+(1-P_{t,l})\bigg{[}1-\prod_{k\in D_{l}}(1-\rho_{kl}P_{t,k})\bigg{]}\right).$
The proof is deferred to the Appendix (Section VI). Next, we proceed directly to Line 1, showing that it may also be derived from the BKF, given $\Pi_{t\mid t-1}=\Pi^{*}(P^{\prime}_{t})$.
Theorem 3
Given a distribution $\Pi_{t\mid t-1}$ such that $\Pi_{t\mid t-1}=\Pi^{*}(P)$ for some $P\in[0,1]^{n}$, we may compute (6) in $\mathcal{O}(n)$ via
$P_{t}=(X\Pi_{t\mid t})_{r}=\bigg{(}X\frac{T_{t}\circ\Pi_{t\mid t-1}}{(T_{t})^{\top}\Pi_{t\mid t-1}}\bigg{)}_{r}=$ (15)
$\frac{py_{t,l}P^{\prime}_{t,l}+(1-y_{t,l})\tilde{p}P^{\prime}_{t,l}}{y_{t,l}[pP^{\prime}_{t,l}+\tilde{q}(1-P^{\prime}_{t,l})]+(1-y_{t,l})[\tilde{p}P^{\prime}_{t,l}+q(1-P^{\prime}_{t,l})]}.$
Again, we defer the proof to the Appendix (Section VI Interestingly, this indicates that the Bayesian update of the observation is equivalent to the BKF’s observation update in this context.
These theorems provide more formal evidence that the idea of using Bayes’ law given the belief $P_{t},P^{\prime}_{t}$ and leveraging the maximum entropy distribution $\Pi^{*}(P)$ is indeed closely related to the actual function of the BKF algorithm. The performance of the MFA algorithm thus depends on how well the BKF function performs when restricted to maximum entropy distributions. The unrestricted variant is capable of computing lower entropy representative distributions, which may potentially lead to better estimates of $\hat{x}_{t}$. This distinction can be regarded as the precise loss of information from storing only $n$ real numbers, in contrast to storing the full distribution over $2^{n}$.
It is important to note that although the MFA approaches can follow each step of the BKF algorithm individually, it is typically true that $\Pi_{t|t-1}\neq\Pi^{*}(X\Pi_{t|t-1})$. This implies that even if the prior $\Pi_{t-1|t-1}=\Pi^{*}(P_{t-1})$ maximized entropy, the assumption of Theorem 3 is not be satisfied and often $P_{t}\neq X\Pi_{t|t}$.
IV-A Simulation Results
We now provide simulations to verify the effectiveness of the MFA algorithm based on the intuition of Theorems 2 and 3. To evaluate the algorithms, we define a True Estimation Rate metric, $\textrm{TER}(x_{t},\hat{x}_{t})=1-(||x_{t}-\hat{x}_{t}||^{2}_{2})/n$, which is the fraction of nodes whose states are correctly estimated. Throughout (unless otherwise stated), the parameters are $\alpha=0.2,p=0.8,q=0.8,\rho_{ij}=0.1$, and we run each experiment for 20 time steps in 100 independent random trials. We assume that the first node is the external threat that is known to be compromised and cannot be cleaned (we also exclude this node for evaluation purposes). The initial state is then $x_{0}=(1,0,0,0,\dots)$; the initial beliefs of the algorithms are $P_{0}=(1,0.5,0.5,\dots)$ and $\Pi_{0}=\Pi^{*}(P_{0})$ for MFA and BKF, respectively. Additionally, at each time step, we assume 2 nodes are randomly selected for cleaning (excluding node 1).
Figure 1: Performance Comparison between the MFA and BKF Estimators. On the top-left we observe a comparative performance between the algorithms on the Computer Network Graph, and in the bottom left we compare runtimes on ring graphs. In the center plots, we observe how the performance on ring graphs varies with the size of network. In the right plots, we report the same experiment but on Erdős-Rényi graphs.
To compare the performance between the MFA and BKF algorithms in terms of TER and runtime, we provide six experiments shown in Figure 1. In the top-left plot, we evaluate the True Estimation Rate at each timestep between both algorithms via an 11-node network proposed in [19], which we term the Computer Network Graph. We take $p=0.2,q=1$ to match that work as closely as possible, noting that all of our nodes would be considered OR nodes (we did not implement their notion of AND nodes in this work) and that we include the external attacker as a node (making it an 11-node graph). We observe that MFA follows BKF relatively closely at all time steps; given that the BKF algorithm is optimal in expectation, it is possible for the MFA algorithm to occasionally outperform BKF on specific sample paths.
The lower-left plot of Figure 1 is a run-time comparison between the two algorithms on ring graphs. In the plot, $n$ denotes the number of nodes and the average runtime of 5 independent simulation runs is recorded for 20 time steps each. As expected, the BKF algorithm is known to take exponential time (and we were able to compute up to $n=14$ for BKF), which is verified by the exponential curve in the plot. It should be noted that for systems not much larger than the ones considered (e.g., a network with 30 nodes), the BKF estimator will no longer be practical, while the MFA estimator is.
The two middle plots present the changes in average performance on ring graphs of variable size, where TER is averaged across both 100 trials and all 20 time steps. In the top-center plot, we evaluate $n\in\{5,\dots,10\}$ and observe that MFA and BKF perform very closely. In the bottom-center plot, we forgo computing BKF and plot MFA’s performance on $n\in\{5,\dots,300\}$. Interestingly, we observe a dramatic increase in performance as the number of nodes increases on small ring graphs, then leveling out around 0.93 for graphs of sufficient size.
In the two rightmost plots, we present the same study as the middle plots, but on Erdős-Rényi (ER) graphs. For these experiments, we define ER graphs where each edge exists with probability $0.2$, and we generate a new graph for each trial. In the top plot, we observe that MFA and BKF follow each other closely; in the bottom plot, MFA’s performance is the strongest on small graphs (which may be easy to predict because they have few edges). Then, we observe a small increase in performance for certain moderately sized graphs, before the performance drops again. Finally, as graphs grow large TER steadily improves again. This suggests that certain edge densities may be more favorable for estimation than others, due to the fixed edge probability across all ER graph sizes.
V Conclusion
We have presented a tractable algorithm for state estimation in a networked FlipIt model. Our theoretical and simulation results show that our proposed method approximates the proven optimal BKF approach well. One interesting open problem is to prove bounds on the accuracy of this estimator.